How many kilometers would you have to go above the surface of the earth for your weight to decrease to 36 % of what it was at the surface?
Answer #1

use,
gh = g*(1-(2*h)/R)
given,
gh = 0.36*g
so,
0.36*g = g*{1-(2*h)/R}
0.36 = 1-(2*h)/R
(2*h)/R = 1-0.36
(2*h)/R = 0.64
2*h = 0.64*R
h = 0.32*R

radius of earth = 6378.1 km
so,
h = 0.32*6378.1 km
= 2040.99 km
= 2041 km

Answer: 2041 km